# High-fidelity Metalens Imaging via Aberration Correcting Transformers

### v1

@@ B.3 Video Aberration Correction
3D-GS [41] 17.09 0.3270 0.8892
3D-GS w/ Ours 25.65 0.6406 0.4244
3D-GS + Ours 25.66 0.6395 0.4201
Table 5: Quantitative results on the aberrated Mip-NeRF360 [6] dataset.
B.5 More ablation studies
We conducted further ablation studies to demonstrate the effectiveness of the proposed methods. We evaluated all models after training 15K iterations.
Comparison on the Number of Filters for MAFG. Tab. 6 shows the results of comparing different numbers of involved filters in MAFG. The more filters we employed, the aberration correction performance was improved as it could enrich the input of the restoration network more. However, the performance dropped when we leveraged 4 filters in MAFG. We assume this is because excessive usage of filters may introduce over-redundancies and conflicting correlations within the representations, hampering correcting aberrations. Thus, we harnessed 3 filters in our experiments.
@@ B.5 More ablation studies
0 26.14 0.7231 0.3296
1 30.73 0.8065 0.2890
2 31.55 0.8108 0.2835
3 32.01 0.8121 0.2852
4 31.92 0.8111 0.2804
Table 6: Ablation study on the number of filters for MAFG.
Ablation on Using Intensity for MAFG. We conducted an ablation study on using channel intensity information when defining filters for MAFG. We obtained multiple adaptive filters following Eq. 3 (MAFG) and simple multiple filters using Eq. 2 (MAFG w/o Intensity) in the main paper. We ran an experiment on a different test dataset, the aberrated DIV2K [1] dataset, to validate the adaptability of the proposed MAFG to unseen data. Exploiting channel intensity is simple, but it could impart the model flexibility to different images and facilitate aberration correction, especially showing a noticeable improvement in LPIPS as described in Tab. 7. We believe that a more sophisticated approach to incorporating image information, such as image intensity, into the filters can further enhance the restoration process.
@@
Restormer [92] 23.03 0.5872 0.4447
MAFG w/o Intensity 28.17 0.7305 0.4184
MAFG 28.28 0.7357 0.3912
Table 7: Ablation study on exploiting channel intensity in MAFG, evaluated on the aberrated DIV2K [1] dataset.
@@
Figure 14: Additional qualitative results on the aberrated Open Image V7 [48, 7] dataset.
Figure 15: Restored images and noise maps on the aberrated Open Image V7  [48, 7] dataset. Images in the third and fifth columns show the computed noise maps. Noise maps are normalized for better visibility.
@@
Figure 16: Additional qualitative results on the metalens-captured images.

### v2

@@ Appendix B Analysis and Discussion
MAFG w/o CA 33.59 0.8697 0.2044
MAFG w/o CI 33.82 0.8728 0.2074
MAFG 33.87 0.8757 0.2036
Table 4: Ablation study on the MAFG.
@@
3 33.60 0.8725 0.2027
7 33.76 0.8737 0.2040
13 33.87 0.8757 0.2036
Table 5: Ablation study on the number of filters of MAFG.
Effect of STAF. Tab. 6 provides experimental results to demonstrate the competence of STAF. Employing transposed self-attention (TA) alone, similar to Restormer [93] often lacks encoding spatial context, leading to the worst performance among the baselines (first row in Tab. 6). In the case of harnessing both spatial self-attention (SA) and TA, the way of fusing two different features also affects the restoration quality. Sequentially applying SA and TA (second row) where several previous works [12, 15, 81] proposed, still fall short of performance as SA and TA are identically implemented in both encoder and decoder. Meanwhile, the proposed STAF (fourth row) can effectively aggregate the outputs from two different modules considering the roles of encoder and decoder. It further facilitates restoration by enforcing queries of attention modules in decoder blocks focusing more on the fine details and acheieves 0.18 dB improvement in PSNR compared to using original queries (third row), resulting in superior restoration quality.
@@
TA - 33.07 0.8671 0.2062
SA, TA Sequential 33.63 0.8690 0.2165
SA, TA Weighted sum 33.69 0.8714 0.2051
SA, TA Weighted sum 33.87 0.8757 0.2036
Table 6: Ablation study on the STAF.
Appendix C Implementation Details
We trained our model for 100K iterations with a batch size of 8 on 4 RTX A6000 GPUs. We adopted Overlapping Cross Attention (OCA), which partitions query matrix with non-overlapping windows and key and value matrices with overlapping windows, from HAT [14] for our SA module, following X-Restormer [12]. We employed the Gated-Deconv Feed-Forward network from Restormer [93] for our feed-forward network (FFN). Please refer to Restormer and X-Restormer for more details.
For the cross-attention module in the MAFG, we projected the deconolved images to 32-dimensional feature space. The cross-attention uses overlapping windows similar to the SA modules in STAF and adopts two attention heads. In the case of image dataset construction, we resized each image such that the longer dimension (either height or width) was set to 1215 pixels, while the shorter dimension was scaled proportionally to fit the images to the sensor size while preserving the original aspect ratio. We set the sensor noise $n$ in Eq. 1 in the main paper as a sum of Gaussian and Poisson noise. Each noise is sampled from the Gaussian distribution $\mathcal{N}(x,\sigma_{g}^{2})$ and the Poisson distribution $\mathcal{P}(\frac{x}{\sigma_{p}})$, respectively, where $x$ is the image, $\sigma_{g}=1e-5$, and $\sigma_{p}=4e-5$, following Tseng et al. [80].
Appendix D Additional Results
In this section, we provide additional experimental results on each task that were not delivered in the main paper due to the page limit. We also attached videos of the 3D reconstruction and video aberration correction task results.
Evaluation Metrics We assessed the proposed method with various evaluation metrics including Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [84], and Learned Perceptual Image Patch Similarity (LPIPS) [95]. We further validated our methods with non-reference metrics, including Multi-scale Image Quality (MUSIQ) [35], Multi-dimension Attention Network Image Quality Assessment (MANIQA) [87], Hyper-network Image Quality Assessment HyperIQA [76] and Natural Image Quality Evaluator (NIQE) [61] on the image datasets.
D.1 Image Aberration Correction
Fig. 16 shows extra qualitative results on the Open Image V7 [44, 6] dataset with spatially variant aberration. The proposed method can mitigate aberrations in both central and peripheral regions and restore fine details. Tab. 7 shows the evaluations on non-uniform aberration datasets under several non-reference metrics, and the proposed method consistently outperforms diverse baseline models.
We also provide additional results on five different uniform aberration datasets, where we used the PSFs at 0°, 5°, 10°, and 15°, 20°  field angle, respectively. As shown in Fig. 17, Restormer [93] shows visually satisfactory results on themildly aberrated dataset (0° and 5° field angle), revealing the potential of ViT in metalens imaging. However, both Restormer and Restormer with a single Wiener filter (Restormer + SWF) struggle with restoring file details or lose image contrast when the images are harshly aberrated (15° and 20° field angle), highlighting sophisticated approach is required. Since the proposed method can model spatially variant aberration, it can also effectively correct uniform aberrations at various field angles.
[…]
Restormer [93] 48.028 0.243 0.340 7.326
Restormer [93] + SWF 48.028 0.239 0.326 7.295
Ours 50.938 0.267 0.360 6.967
Table 7: Quantitative results on non-reference evaluation metrics.
Uniform aberration at $\theta=5$° Uniform aberration at $\theta=10$° Uniform aberration at $\theta=15$°
[…]
Wiener deconvolution [86] 27.14 0.6653 0.4316 26.72 0.6455 0.4661 26.61 0.6479 0.4656
Eboli et al. [21] 20.75 0.5363 0.6560 20.36 0.5172 0.6909 19.40 0.4794 0.7512
DeblurGANv2 [43] 25.18 0.7286 0.3041 24.51 0.7025 0.3335 23.53 0.6599 0.4036
DWDN [19] 29.04 0.8256 0.2505 27.99 0.8010 0.2691 27.09 0.7763 0.2969
Tseng et al. [80] 29.91 0.8366 0.2513 29.59 0.8213 0.2587 29.14 0.8141 0.2701
Restormer [93] 30.21 0.8654 0.1956 31.51 0.8671 0.2045 30.98 0.8526 0.2202
Restormer [93] + SWF 33.07 0.8672 0.2227 32.10 0.8434 0.2504 31.21 0.8337 0.2630
Ours 34.83 0.8904 0.1795 34.39 0.8788 0.1979 34.05 0.8708 0.2070
Table 8: Quantitative results Open Image V7 [44, 6] dataset with spatially varying and invariant aberration, where $\theta$ denotes a field angle.
D.2 Images from the Fabricated Metalens
Model Fine-tuning Details We fine-tuned our pre-trained model, from Fig. 10 in the main paper, using images captured with the fabricated metalens. The fine-tuning dataset includes 200 images for training and 20 images for validation. However, they appeared to be dark and exhibited significant noise. To mitigate these issues, we implemented a preprocessing pipeline from Shen et al. [70]. Specifically, we applied Contrast-Limited Adaptive Histogram Equalization (CLAHE) [65] to enhance contrast, utilized an improved Anscombe transformation in the wavelet domain, and performed total variation-based denoising before reconstructing the images.
We fine-tuned the pre-trained model for 5,500 iterations and the fine-tuned network could produce promising results, despite the limited dataset size.
Results Fig. 18 presents the qualitative results of images captured with the fabricated metalens and restored ones using the proposed method and other baseline models. All models are fine-tuned under the same experimental settings. The captured images are harshly degraded, but the proposed method could effectively restore the overall color fidelity of the images with fine details, while other baseline models produce noisy and low-contrast outputs. Such outcomes can demonstrate the adaptability of the proposed method to real-world scenarios.
D.3 Video Aberration Correction
Fig. 19 shows additional qualitative results on the aberrated DVD [75] dataset. Similar to Fig. 10 in the main paper, the proposed method could robustly address aberrations in the video and produce aberration corrected representations.
@@ D.3 Video Aberration Correction
Figure 16: Additional qualitative results on the Open Image V7 [44, 6] dataset with spatially variant aberration.
Figure 17: Additional qualitative results on the Open Image V7 [44, 6] dataset with uniform aberration at different field angles.
@@
Figure 18: Additional qualitative results on the captured images with the fabricated metalens.
@@
D.4 3D Reconstruction with Aberrated Images
Implementation Details We set the loss weighting hyperparameter $\lambda=1.0$ for the Tanks&Temples [40] datasets. For the LLFF [60] dataset, we set $\lambda=0.75$ and inactivated periodically resetting opacity of 3D Gaussians to preserve more points in the forward-facing dataset [46]. We use default hyperparameters including learning rate, training iterations and 3D Gaussian densification parameters from the original paper [37] for our experiments.
Results We provide extra qualitative results on the LLFF [60] and Tanks&Temples [40] datasets in Fig. 20 and Fig. 21. We further conducted experiments on Mip-NeRF 360 [5] dataset with spatially varying aberration, where we set $\lambda=1.0$. As described in Tab. 9, the proposed training pipeline that embed our pre-trained restoration network into 3D-GS could improve the scene reconstruction quality through additional multi-view consistency guidance. Fig. 22 shows that 3D-GS could not model the spatially varying aberration of metalens and failed to produce clean representations. On the other hands, the proposed method enables aberration modeling, demonstrating the capability of clean 3D reconstruction from the set of multi-view aberrated images of our approach.
[…]
3D-GS [37] 19.58 0.3857 0.8166
3D-GS w/ Ours 25.88 0.6669 0.3696
3D-GS + Ours 25.94 0.6676 0.3693
Table 9: Quantitative results on the Mip-NeRF 360 [5] dataset with spatially varying aberration.
Figure 20: Qualitative results on the aberrated LLFF [40] dataset.
Figure 21: Qualitative results on the aberrated Tanks&Temples [40] dataset.
@@
Appendix E Limitations and Future Work
We fabricated a metalens using the phase profile of [80] and captured images projected on the display, following its optical setup. However, capturing actual scenes can introduce various physical variables, such as multiple light sources and noises, which can undermine the quality of the images, but the current optical system is not well-suited for in-the-wild imaging. To address this limitation, we are designing a new lens optimized for real world applications, and propose a more robust and effective restoration network capable of handling metalens-captured scenes as a future work.